### v1

# Trajectory Optimization Under Stochastic Dynamics Leveraging Maximum Mean Discrepancy

Abstract
This paper addresses sampling-based trajectory optimization for risk-aware navigation under stochastic dynamics. Typically such approaches operate by computing $\tilde{N}$ perturbed rollouts around the nominal dynamics to estimate the collision risk associated with a sequence of control commands. We consider a setting where it is expensive to estimate risk using perturbed rollouts, for example, due to expensive collision-checks. We put forward two key contributions. First, we develop an algorithm that distills the statistical information from a larger set of rollouts to a reduced-set with sample size $N<<\tilde{N}$. Consequently, we estimate collision risk using just $N$ rollouts instead of $\tilde{N}$. Second, we formulate a novel surrogate for the collision risk that can leverage the distilled statistical information contained in the reduced-set. We formalize both algorithmic contributions using distribution embedding in Reproducing Kernel Hilbert Space (RKHS) and Maximum Mean Discrepancy (MMD). We perform extensive benchmarking to demonstrate that our MMD-based approach leads to safer trajectories at low sample regime than existing baselines using Conditional Value-at Risk (CVaR) based collision risk estimate.
I Introduction
Risk-aware trajectory optimization provides a rigorous template for assuring safety under stochastic dynamics model [1], [2]. In this paper, we adopt the premise of sampling-based optimization that relies on simulating the stochastic dynamics [1], [3]. These class of approaches can be applied to arbitrary noise model and even non-differentiable black-box functions. As a result, they allow us to relax the linear dynamics and Gaussian noise assumption prevalent in many existing works, e.g, [4], [5].
Figure 1: A standard pipeline for risk-aware optimization based on control sampling along with our improvement. These class of approaches rely on simulating the forward dynamics of the vehicle to obtain $\tilde{N}$ samples from the state trajectory distribution, which are then used to estimate risk. Our work provides a novel risk-surrogate and a systematic way of estimating it using a reduced number ($N$) of state trajectory samples (a.k.a the reduced-set).
A typical pipeline is presented in Fig.1. Given a sequence of control commands, we compute $\tilde{N}$ forward simulations a.k.a rollouts of the stochastic dynamics, resulting in as many samples of state trajectories. We then evaluate the state-dependent cost/constraints (e.g safety distance violations) along the state trajectory samples. This is followed by computing some risk-aware statistics such as Conditional Value at Risk (CVaR) of the cost/constraint samples.
In this paper, we consider a setting, where cost/constraint evaluations along the state trajectory samples are computationally expensive, e.g due to expensive collision checks. Furthermore, often, computing the rollouts itself could be computationally and memory intensive. Our key idea is to distill the statistical information contained in $\tilde{N}$ state trajectory samples into a reduced-set with sample size $N<<\tilde{N}$. For example, in our implementation, $\tilde{N}\approx N^{2}$. Furthermore, the distillation to reduced-set should be done in a way that it ensures reliable downstream risk estimation using only the $N$ samples.
Algorithmic Contribution: We formalize our key ideas using distribution embedding in Reproducing Kernel Hilbert Space (RKHS) and the associated Maximum Mean Discrepancy (MMD) measure [6],[7]. Specifically, we formulate a MMD-based surrogate for state dependent risk. The important feature of our risk surrogate is that we can apply the tools from RKHS distribution embedding to improve its sample efficiency. Put differently, we propose a bi-level optimization that provides a systematic way of reducing the number of state trajectory samples needed to estimate the risk associated with a sequence of control commands. Finally, we present a custom sampling-based optimizer to minimize the proposed risk surrogate.
State-of-the-Art Performance: We perform two sets of benchmarking to demonstrate the advantages of our approach. First, we show that our MMD-based surrogate is more sample efficient in estimating collision risk in a given scene, as compared to popular alternatives based on CVaR [1],[8],[9],[10],[11]. The difference is particularly stark in the low sample regime. Second, we also apply our proposed trajectory optimizer in a Model Predictive Control(MPC) setting in a high-fidelity simulator CARLA[12] to highlight the improvements over a deterministic noise-ignorant baseline as well as a CVaR-based baseline.
II Problem Formulation
Symbols and Notations
We use small/upper case normal-font to represent scalars. The bold-face small fonts represent vectors while upper-case variants represent matrices. We use $p_{(.)}$ to denote probability density of (.)(.)( . ) and $P$ to represent probability. We use $\langle.,.\rangle_{\mathcal{H}}$ to represent the inner prodct in RKHS $\mathcal{H}$.
II-A Trajectory Optimization
We formulate trajectory planning of the ego-vehicle in the road-aligned Frenet frame, described with the help of a reference path (or road center-line). The variable $s$ and $d$ will represent the longitudinal and lateral displacement in the Frenet Frame, while $\psi$ will represent the heading of the vehicle with respect to the reference path. With these notations in place, we define the risk-aware stochastic trajectory optimization in the following manner.
$\min_{\mathbf{a},\boldsymbol{\theta}}w_{1}E[c(\mathbf{x})]+w_{2}r(\mathbf{x})+w_{3}\left\|\begin{matrix}\mathbf{a}\\ \boldsymbol{\theta}\end{matrix}\right\|_{2}^{2},$ (1a)
$\mathbf{x}_{k+1}=f(\mathbf{x}_{k},a_{k}+\epsilon_{a,k},\theta_{k}+\epsilon_{\theta,k}),\qquad\mathbf{x}_{0}\sim p_{0},$ (1b)
$\theta_{min}\leq\theta_{k}\leq\theta_{max},a_{min}\leq a_{k}\leq a_{max}\forall k,$ (1c)
where $\mathbf{x}_{k}=(s_{k},d_{k},\psi_{k},\dot{\psi}_{k},v_{k})$ represents the state of the vehicle at time-step $k$. The vector $\mathbf{x}$ is the concatenation of the states at different $k$. The function $f$ is taken from [13]. The variable $v_{k}$ is the longitudinal velocity of the ego-vehicle. The control inputs are the longitudinal acceleration $a_{k}$ and steering input $\theta_{k}$. The vectors $\mathbf{a}$ and $\boldsymbol{\theta}$ are formed by stacking $a_{k}$ and $\theta_{k}$ at different time-step $k$ respectively. The stochastic disturbances acting on the vehicle are modeled as an effect of the perturbation of the nominal acceleration and steering inputs by $\epsilon_{a,k}$ and $\epsilon_{\theta,k}$ respectively. Let $\boldsymbol{\epsilon}_{a}$, $\boldsymbol{\epsilon}_{\theta}$ be vectors formed by stacking $\epsilon_{a,k}$ and $\epsilon_{\theta,k}$ respectively at different $k$. We assume that $\boldsymbol{\epsilon}_{a}\sim p_{a}$, $\boldsymbol{\epsilon}_{\theta}\sim p_{\theta}$. The perturbation at the control level is mapped to the state trajectory distribution $p_{\mathbf{x}}$ through $f$. For the sake of generality, we assume that $p_{a}$ and $p_{\theta}$ are dependent on the control inputs. For example, the slippage of a vehicle on icy-roads depends on the magnitude of the acceleration and steering commands. We don’t make any assumptions on the parametric form of $p_{a},p_{\theta}$ and $p_{\mathbf{x}}$. Instead, we just rely on the ability to sample from $p_{a},p_{\theta}$ and rollout the dynamics for every sampled ${\epsilon}_{a,k},{\epsilon}_{\theta,k}$.
The first term in the cost function (1a) minimizes the expected cost over the state distribution. Typically, these would contain costs such as path-following error for which average performance is satisfactory. In contrast, the second term captures the risk associated with state trajectory distribution $\mathbf{x}$ and depends on its higher-order noise characteristics like variance, skewness, kurtosis, etc. Intuitively, the risk captures the probability of a occurrence of an event (e.g collision) for a given sequence of control inputs. We present the algebraic form our risk model in the next subsection. The last terms in (1a) penalizes the use of large control inputs. Each term in the cost function is weighted by $w_{i}$ that can be used to tune the risk-seeking behavior of the ego-vehicle. We also consider control bounds through inequalities (1c).
II-B Algebraic Form of Risk
Let $h_{k}(\mathbf{x}_{k})\leq 0,\forall k$ represent state dependent safety constraints. We can eliminate the temporal dependency by defining the worst-case constraint as $h(\mathbf{x})=\max_{k}(h_{k})$. Clearly, $h(\mathbf{x})\leq 0$ implies $h_{k}\leq 0,\forall k$.
In the stochastic setting where $\mathbf{x}_{k}$ (or $\mathbf{x}$) is a random variable, constraint satisfaction is more appropriately described in terms of so-called chance constraints. These have the general form of $P(h(\mathbf{x})\geq 0)\leq\varepsilon$, where $P$ represents probability and $\varepsilon$ is some constant. Thus, we define risk as
$r(\mathbf{x})=P(h(\mathbf{x})\geq 0)$ (2)
Intuitively, our risk model captures the probability of safety constraints being violated for a given distribution of disturbances. Thus, minimizing it within (1a)-(1c) will lead to safer trajectories. For arbitrary $p_{a},p_{\theta}$ and/or highly-non linear dynamics model $f$, constraint function $h$, the analytical form of r.h.s of (2) is often not known. Thus, existing literature on risk-aware trajectory optimization [1],[8], [14] focuses on developing computationally tractable surrogates.
Figure 2: The mass of $p_{\overline{h}}$ is to the right of $\overline{h}=0$. The optimal control input is one that leads to state-trajectory distribution for which $p_{\overline{h}}$ resembles a Dirac-Delta distribution.
III Main Algorithmic Results
Our approach follows the typical pipeline shown in Fig.1. The unique feature is our risk surrogate and how it can be estimated using only a subset of state-trajectory rollouts. We present the key building blocks next.
III-A State Risk as Difference of Distributions
Let us define a constraint residual function as
$\overline{h}(\mathbf{x})=\max(0,h(\mathbf{x})).$ (3)
In the deterministic scenario, driving $\overline{h}(\mathbf{x})$ to zero will push $h(\mathbf{x})$ to the feasible boundary. In the stochastic case, (3) maps $p_{\mathbf{x}}$ to a distribution of constraint residuals. Let $\overline{h}(\mathbf{x})\sim p_{\overline{h}}$. The key insight in our work is that although we don’t know the parameteric form for $p_{\overline{h}}$, we can be certain that its entire mass lies to the right of $\overline{h}=0$ (see Fig.2). Moreover, as $P(h(\mathbf{x})\geq 0)$ approaches zero, the $p_{\overline{h}}$ converges to a Dirac-Delta distribution $p_{\delta}$. In other words, one way of reducing risk is to minimize the difference between $p_{\overline{h}}$ and $p_{\delta}$. Thus, we propose the following risk estimate following our prior works [15], [16].
$r(\mathbf{x})\approx\mathcal{L}_{dist}(p_{\delta},p_{\mathbf{x}}),$ (4)
where $\mathcal{L}_{dist}$ is any measure that characterizes the difference between two distributions. In the subsequent sections, we propose MMD as a potential choice for $\mathcal{L}_{dist}$.
III-B RKHS Embedding of Functions of Random Variables
Our approach builds on the ability of embedding functions of random variables in the RKHS [6], [17]. For example, state trajectory $\mathbf{x}$ is a function of two random variables $\boldsymbol{\epsilon}_{a}$, $\boldsymbol{\epsilon}_{\theta}$, formed by stacking $\epsilon_{a,k}$ and $\epsilon_{\theta,k}$ at different time-step $k$. The RKHS embedding of $\mathbf{x}$ denoted by $\mu[\mathbf{x}]$ is computed as
$\mu[\mathbf{x}]=E[\phi(\mathbf{x}(\boldsymbol{\epsilon}_{a},\boldsymbol{\epsilon}_{\theta}))]=E[K_{\sigma}(\mathbf{x}(\boldsymbol{\epsilon}_{a},\boldsymbol{\epsilon}_{\theta}),.)],$ (5)
where $E[.]$ stands for the expectation operator and $\phi$ is a non-linear transformation commonly referred to as the feature-map [17]. One of the key properties of $\phi$ is that the inner product in the RKHS $\left\langle\phi(\mathbf{z}),\phi(\mathbf{z}^{^{\prime}})\right\rangle_{\mathcal{H}}$ can be expressed as $K_{\sigma}(\mathbf{z},\mathbf{z}^{^{\prime}})$ for any arbitrary vector $\mathbf{z},\mathbf{z}^{^{\prime}}$. Here,$K_{\sigma}$ is a positive definite function known as the kernel function with hyper-parameter $\sigma$. Throughout this paper, we used the Laplacian kernel for which $\sigma$ represents the kernel-width.
Typically, the r.h.s of (5) is difficult to compute. Thus, it is common to compute the empirical estimate by replacing the expectation operator through sample mean in the following manner.
$\hat{\mu}[\mathbf{x}]=\sum_{i,j=1}^{i,j=N}\frac{1}{N^{2}}\phi({{}^{ij}}\mathbf{x})=\sum_{i,j=1}^{i,j=N}\frac{1}{N^{2}}K_{\sigma}({{}^{ij}}\mathbf{x},.),$ (6)
where, ${{}^{ij}}\mathbf{x}=\mathbf{x}({{}^{i}}\boldsymbol{\epsilon}_{a},{{}^{j}}\boldsymbol{\epsilon}_{\theta})$ and ${{}^{i}}\boldsymbol{\epsilon}_{a}$, ${{}^{j}}\boldsymbol{\epsilon}_{\theta}$ are i.i.d samples of $\boldsymbol{\epsilon}_{a}$, $\boldsymbol{\epsilon}_{\theta}$ respectively. Following a similar approach, the RKHS embedding of $\overline{h}(\mathbf{x})$ (or $p_{\overline{h}}$) (function of random variable $\mathbf{x}$), and its empirical estimate can be computed as (7a)-(7b), wherein, ${{}^{ij}}\overline{h}=\overline{h}({{}^{ij}}\mathbf{x})$
$\mu[\overline{h}]=E[\phi(\overline{h}(\mathbf{x}))],$ (7a)
$\hat{\mu}[\overline{h}]=\sum_{i,j=1}^{i,j=N}\frac{1}{N^{2}}\phi({{}^{ij}}\overline{h})=\sum_{i,j=1}^{i,j=N}\frac{1}{N^{2}}K({{}^{ij}}\overline{h},.),$ (7b)
III-C MMD Based Risk Surrogate
Let $\delta$ be a random variable with Dirac-Delta distribution. Let $\mu[\delta]$ be the RKHS embedding of $\delta$ (or $p_{\delta}$). We use the Maximum Mean Discrepancy (MMD) between $p_{\overline{h}}$ and $p_{\delta}$ as our choice for $\mathcal{L}_{dist}$ in (4)
$r_{MMD}=\mathcal{L}_{dist}(p_{\delta},p_{\mathbf{x}})=\overbrace{\left\|\mu[\overline{h}(\mathbf{x})]-\mu[\delta]\right\|_{\mathcal{H}}^{2}}^{MMD}.$ (8)
It can be shown that $r_{MMD}=0$ implies $p_{\overline{h}}=p_{\delta}$ [6], [17], [7]. In other words, $r_{MMD}=0$ implies that a state trajectory is safe with probability one. From practical stand-point, since we will only have access to state trajectory samples obtained by roll-out of the dynamics, we have to resort to the empirical (biased) MMD estimate [7] given by
$r_{MMD}^{emp}=\left\|\hat{\mu}[\overline{h}(\mathbf{x})]-\hat{\mu}[\delta]\right\|_{\mathcal{H}}^{2},$ (9)
where, $\hat{\mu}[\delta]$ is computed based on the $N^{2}$ samples of $\delta$ drawn from $p_{\delta}$ [^1] We can approximate $p_{\delta}$ through a Gaussian $\mathcal{N}(0,\epsilon)$, with an extremely small covariance $\epsilon$ ($\approx 10^{-5}$).. The empirical $r_{MMD}^{emp}$ converges to $r_{MMD}$ at the rate proportional to $\frac{1}{\sqrt{N}}$ [17], [7].
Why is $r_{MMD}^{emp}$ an Ideal Choice for Risk Estimate? Eqn.(7b) suggests that we need $N^{2}$ rollouts of state trajectory $\mathbf{x}$ and evaluate $\overline{h}(\mathbf{x})$ over each of them to compute $\hat{\mu}[\overline{h}]$ and consequently $r_{MMD}^{emp}$. However, this can be computationally prohibitive if evaluating $\overline{h}$ is difficult, e.g, due to expensive collision checks. Fortunately, RKHS embedding and MMD provide us a set of tools to systematically choose only $N$ out of those $N^{2}$ rollouts to estimate $\overline{h}(\mathbf{x})$ and $r_{MMD}^{emp}$ and yet ensure minimal loss in risk estimation accuracy.
Let ${{}^{l}}\mathbf{x}^{^{\prime}},l=1,2,\dots,N$ be some $N$ subset/reduced-set of $N^{2}$ samples of ${{}^{ij}}\mathbf{x}$. Moreover, let us re-weight the importance of each ${{}^{l}}\mathbf{x}^{^{\prime}}$ through ${{}^{l}}\beta$ such that $\sum_{l}{{}^{l}}\beta=1$. Then, the RKHS embeddings of the state trajectory and constraint residual distribution using the ${{}^{l}}\mathbf{x}^{^{\prime}}$ samples are given by
$\hat{\mu}[\mathbf{x}^{^{\prime}}]=\sum_{l=1}^{l=N}{{}^{l}}\beta\phi({{}^{l}}\mathbf{x}^{^{\prime}}),\qquad\hat{\mu}[\overline{h}^{^{\prime}}]=\sum_{l=1}^{l=N}{{}^{l}}\beta\phi(\overline{h}({{}^{l}}\mathbf{x}^{^{\prime}}))$ (10)
Now, Theorem 1 from [6] ensures that if $\hat{\mu}[\mathbf{x}^{^{\prime}}]$ is close to $\hat{\mu}[\mathbf{x}]$ in MMD sense, then $\hat{\mu}[\overline{h}^{^{\prime}}]$ will be close to $\hat{\mu}[\overline{h}]$ in the same metric. Consequently, $r_{MMD}^{emp}$ computed on the smaller $N$ rollouts will be close to that computed over the larger $N^{2}$ samples. Alternately, if we can minimize $\left\|\hat{\mu}[\mathbf{x}^{^{\prime}}]-\hat{\mu}[\mathbf{x}]\right\|_{\mathcal{H}}^{2}$ , then we can ensure minimal loss in risk estimation accuracy while reducing the sample-size. This is discussed in the next subsection.
III-D Optimal Reduced Set
There are three ways in which we can minimize $\left\|\hat{\mu}[\mathbf{x}^{^{\prime}}]-\hat{\mu}[\mathbf{x}]\right\|_{\mathcal{H}}^{2}$. We can choose the optimal subset/reduced-set out of $N^{2}$ dynamics rollouts. Moreover, we can optimize ${{}^{l}}\beta$ as well as the Kernel parameter $\sigma$. To this end, we formulate the following bi-level optimization, wherein $\mathbf{O}$ is a matrix formed by row-wise stacking of the $N^{2}$ samples of $\mathbf{x}$.
$\min_{\boldsymbol{\lambda},\sigma}$ $\left\|\sum_{i,j=1}^{i,j=N}\frac{1}{N^{2}}\phi({{}^{ij}}\mathbf{x})-\sum_{l=1}^{l={N}}{{}^{l}}\beta^{*}\phi({{}^{l}}\mathbf{x}^{^{\prime}})\right\|_{\mathcal{H}}^{2}$ (11a)
$F_{\boldsymbol{\lambda}}({\mathbf{O}})=({{}^{1}}\mathbf{x}^{^{\prime}},{{}^{2}}\mathbf{x}^{^{\prime}},\dots,{{}^{N}}\mathbf{x}^{^{\prime}})$ (11b)
${{}^{l}}\beta^{*}=\begin{aligned} &&&\arg\min_{{{}^{l}}\beta}\left\|\sum_{i,j=1}^{i,j=N}\frac{1}{N^{2}}\phi({{}^{ij}}\mathbf{x})-\sum_{l=1}^{l={N}}{{}^{l}}\beta\phi({{}^{l}}\mathbf{x}^{^{\prime}})\right\|_{\mathcal{H}}^{2}\\ &&&\operatorname{s.t.}\sum_{l}{{{}^{l}}\beta}=1\end{aligned}$ (11c)
In the above optimization, ${F}_{\boldsymbol{\lambda}}$ is a function that chooses $N$ rows out of $\mathbf{O}$ to provide the reduced-set samples. It is parameterized by vector $\boldsymbol{\lambda}$. That is, different choices of $\boldsymbol{\lambda}$ lead to different reduced-set selection. Note that the cost terms (11a), (LABEL:bi_level_inner_red_cost) can be computed via the kernel trick [6].
As can be seen, the inner optimization (LABEL:bi_level_inner_red_cost) is defined over just the weights ${{}^{l}}\beta$ for a fixed reduced-set selection given by $\boldsymbol{\lambda}$. The outer optimization in turn optimizes in the space of $\boldsymbol{\lambda}$ and kernel parameter in order to reduce the MMD cost associated with optimal ${{}^{l}}\beta^{*}$. We use the approach presented in [18] that combines gradient-free cross entropy method (CEM) [19] with quadratic programming (QP) to solve (11a)-(LABEL:bi_level_inner_red_cost). Specifically, we sample $\boldsymbol{\lambda},\sigma$ from a Gaussian distribution and solve the inner optimization for each of the samples to obtain ${{}^{l}}\beta^{*}$. We then evaluate the upper-level cost for all $({{}^{l}}\beta^{*},\boldsymbol{\lambda},\sigma)$ and subsequently modify the sampling distribution to push down this cost. We leverage the fact that the inner optimization (LABEL:bi_level_inner_red_cost) is essentially a equality constrained QP with a closed-form solution to develop a heavily parallelized solver over GPUs.
Selection Function: Let $\boldsymbol{\lambda}\in\mathbb{R}^{N^{2}}$ be an arbitrary vector and $\lambda_{t}$ be its $t^{th}$ element. Assume that $|\lambda_{t}|$ encodes the value of choosing the sample of the state trajectory $\mathbf{x}$ stored in the $t^{th}$ row of ${\mathbf{O}}$. That is, larger the $|\lambda_{t}|$, the higher the impact of choosing the $t^{th}$ row of ${\mathbf{O}}$ in minimizing (11a). With these constructions in place, we can now define $F_{\boldsymbol{\lambda}}$ through the following sequence of mathematical operations.
$\mathcal{S}=Argsort\{|\lambda_{1}|,|\lambda_{2}|,\dots|\lambda_{N^{2}}|)\}$ (12a)
$\mathbf{O}^{{}^{\prime}}=[{\mathbf{O}}_{t_{1}},{\mathbf{O}}_{t_{2}},\dots,{\mathbf{O}}_{t_{N^{2}}}],\forall t_{e}\in\mathcal{S}$ (12b)
$F_{\boldsymbol{\lambda}}({\mathbf{O}})=\mathbf{O}^{{}^{\prime}}_{t_{N^{2}-N}:t_{N^{2}}}=({{}^{1}}\mathbf{x}^{^{\prime}},{{}^{2}}\mathbf{x}^{^{\prime}},\dots,{{}^{N}}\mathbf{x}^{^{\prime}})$ (12c)
As can be seen, we first sort in increasing value of $|\lambda_{t}|$ and compute the required indices $t_{e}$. We then use the same indices to shuffle the rows of $\mathbf{O}$ in (12b) to form an intermediate variable $\mathbf{O}^{{}^{\prime}}$. Finally, we choose the last $N$ rows of $\mathbf{O}^{{}^{\prime}}$ as our reduced-set in (12c). Intuitively, (12a)-(12c) parameterizes the sub-selection of the rows of ${\mathbf{O}}$ to form the reduced-set. That is, different $\boldsymbol{\lambda}$ gives us different reduced-sets. Thus, a large part of solving (11a)-(LABEL:bi_level_inner_red_cost) boils down to arriving at the right $\boldsymbol{\lambda}$. As mentioned above, this is done through a combination of gradient-free search and quadratic programming.
III-E Trajectory Optimizer
Alg.1 presents our approach for solving (1a)-(1c) when the risk cost is given by $r_{MMD}^{emp}$. Our approach combines concepts from constrained version of gradient-free Cross Entropy Method (CEM) [18] and, Model Predictive Path Integral (MPPI) [20] and convex optimization. The basic idea is to use sampling to iteratively arrive at control inputs that have low risk and expected cost as well as small magnitudes.
Alg.1 begins by initializing the mean and covariance for sampling in Line 2. However, instead of directly sampling control inputs, we sample longitudinal velocity and lateral-offsets (from reference path) set-points from a Gaussian distribution (Line 5). This is then fed to a Frenet planner (Line 6) inspired from [21] that generates a batch of trajectories. These trajectories are then converted to acceleration and control inputs by using the differential flatness property of a car-like vehicle [22]. We also clip the magnitudes to satisfy the control bounds. In Line 7, we sample the control perturbation and perform $N^{2}$ rollouts of discrete-time dynamics (1b) to obtain as many state trajectory samples for every control input sample. Note that this rollout process is repeated for every sample in the batch of control input generated in line 6. In line 8, we invoke the optimal reduced-set optimization to choose $N$ subset out of the $N^{2}$ state trajectory samples. In line 9, we estimate $r_{MMD}^{emp}$ over the reduced-set samples. In line 10, we form a $ConstraintElliteSet$ that stores the $n_{c}$ control inputs and Frenet parameters from the total batch that resulted in lowest $r_{MMD}^{emp}$. In line 11, we evaluate the risk, expected and control cost over the control samples and state trajectories stored in $ConstraintElliteSet$. In line 12, we append the costs computed over $ConstraintElliteSet$ to a list. In line 14, we form $ElliteSet$ by choosing $n_{e}$ samples of control and state trajectory from $ConstraintElliteSet$. In line 15, we update the sampling distribution of Frenet parameters through (13a)-(13b), using the samples from the $ElliteSet$.
${{}^{m+1}}\boldsymbol{\nu}=(1-\eta){{}^{m}}\boldsymbol{\nu}+\eta\frac{\sum_{q=1}^{q=n_{e}}t_{q}\mathbf{b}_{q}}{\sum_{q=1}^{q=n_{e}}t_{q}},$ (13a)
${{}^{m+1}}\boldsymbol{\Sigma}=(1-\eta){{}^{m}}\boldsymbol{\Sigma}+\eta\frac{\sum_{q=1}^{q=n_{e}}t_{q}(\mathbf{b}_{q}-{{}^{m+1}}\boldsymbol{\nu})(\mathbf{b}_{q}-{{}^{m+1}}\boldsymbol{\nu})^{T}}{\sum_{q=1}^{q=n_{e}}t_{q}}$ (13b)
$t_{q}=\exp{\frac{-1}{\gamma}(c_{q}})$ (13c)
In (13c), $\gamma$ is the so-called temperature parameter of MPPI [20]. The constant $\eta$ is the learning-rate to prevent sharp changes in the mean and covariance of the sampling distribution. One advantage of Alg.1 is that we have parametrized control sampling over a long horizon in terms of low dimensional vectors of velocity and lateral offset-setpoints. This in turn improves the computational efficiency of Alg.1
1 $M$ = Maximum number of iterations
2 Initiate mean ${}^{m}\boldsymbol{\nu},^{m}\boldsymbol{\Sigma}$, at iteration $m=0$ for sampling frenet parameters (velocity and lane-offsets) $\mathbf{b}$
3 for $m=1,m\leq M,m++$ do
4       Initialize $CostList$ = []
5
6      Draw ${n}$ samples $(\mathbf{b}_{1},\mathbf{b}_{2},\mathbf{b}_{q},....,\mathbf{b}_{{n}})$ from $\mathcal{N}(^{m}\boldsymbol{\nu},^{m}\boldsymbol{\Sigma})$
7
8      Query Frenet Planner $\forall\mathbf{b}_{q}$ : $(\mathbf{a}_{q},\boldsymbol{\theta}_{q})=\text{Frenet Planner}(\mathbf{b}_{q}),\forall q=(1,2,\dots,n)$
9
10      Compute $N$ samples each of $\boldsymbol{\epsilon}_{a},\boldsymbol{\epsilon}_{\theta}$ and subsequently $N^{2}$ rollouts ${{}^{ij}}\mathbf{x}_{q}$ for $(\mathbf{a}_{q},\boldsymbol{\theta}_{q})$ control trajectory. Repeat this process $\forall q=(1,2,\dots,n)$
11
12      Choose $N$ rollouts ${{}^{l}}\mathbf{x}^{^{\prime}}_{q}$ out of $N^{2}$ ${{}^{ij}}\mathbf{x}_{q}$ through (11a)-(LABEL:bi_level_inner_red_cost) and compute corresponding ${{}^{l}}\beta_{q}$ and kernel parameter $\sigma_{q}$. Repeat this process $\forall q=(1,2,\dots,n)$
13
14      Compute $\hat{\mu}[\overline{h}^{^{\prime}}]$ over the optimal reduced-set through (10) and subsequently $r_{MMD}^{emp}$. Repeat this $\forall q=(1,2,\dots,n)$
15
16      $ConstraintEliteSet\leftarrow$ Select top $n_{c}$ batch of $\mathbf{a}_{q},\boldsymbol{\theta}_{q}$, ${{}^{ij}\mathbf{x}_{q}}$, $\mathbf{b}_{q}$ with lowest $r_{MMD}^{emp}$
17
18      Define $c_{q}=w_{1}\sum_{i,j=1}^{i,j=N}c({{}^{ij}}\mathbf{x}_{q})+w_{2}r_{MMD}^{emp}+w_{3}\left\|\begin{matrix}\mathbf{a}_{q}\\ \boldsymbol{\theta}_{q}\end{matrix}\right\|_{2}^{2}$
19       $cost\leftarrow$ $c_{q}$, $\forall q$ in the $ConstraintEliteSet$
20
21      append each computed ${cost}$ to $CostList$
22
23      $EliteSet\leftarrow$ Select top $n_{e}$ samples of ($\mathbf{a}_{q},\boldsymbol{\theta}_{q}$), ${{}^{ij}\mathbf{x}_{q}}$, $\mathbf{b}_{q}$ with lowest cost from $CostList$.
24       $({{}^{m+1}}\boldsymbol{\nu},{{}^{m+1}}\boldsymbol{\Sigma})\leftarrow$ Update distribution based on $EliteSet$
25 end for
return Control Inputs $\mathbf{a}_{q}$ and $\boldsymbol{\theta}_{q}$ corresponding to lowest cost in the $EliteSet$
Algorithm 1 Sampling-Based Optimizer to Solve (1a)-(1c)
IV Connections to Existing Works
Linear Dynamics and Constraints Under Gaussian Noise: For linear dynamics perturbed by Gaussian noise and affine per-step constraint function $h_{k}$, the risk as defined in (2) has an exact convex reformulation [5]. This ensures that the resulting trajectory optimization can be efficiently solved. For non-linear systems and constraint functions, one can resort to linearization to achieve similar structure [4], [23]. However, the state and constraint distribution resulting from linearized system may be inaccurate and compromise the resulting safety guarantees. Recently, authors in [24] presented tractable reformulations for $r$ under Gaussian Mixture Model. Our work differs from all these cited works in the sense that it puts no assumptions on the uncertainty, constraints or dynamics. In fact, our approach can work when all these system specifications are black-box.
Arbitrary Dynamics, Constraints and Noise Model: Our proposed work falls in this category. Typically, when dealing with non-linear dynamics and constraint functions in their exact form (without linearization), under arbitrary noise, the risk $r$ does not have an analytical form. Thus, existing works focus on approximating risk through samples drawn from the state distribution and evaluating the constraint function on them. For example, [8], [9],[10],[11] approximate $r$ through Conditional Value at Risk (CVaR) estimate over the samples of $h({{}^{ij}}\mathbf{x})$. Authors in [1] follow a similar approach but define $r$ as CVaR over the constraint residual function $\overline{h}({{}^{ij}}\mathbf{x})$. Another possibility is to represent $r$ through the so-called sample average approximation (SAA) of $\overline{h}({{}^{ij}}\mathbf{x})$ samples [14].
A key differentiating factor between the existing CVaR [1], [8], [9],[10],[11], SAA [14] based approximations and our MMD-based risk surrogate, is that we can maximize the expressive capacity of the latter for a given sample size (recall (11a)-(LABEL:bi_level_inner_red_cost)). We are not aware of any such mechanism being presented in existing works for CVaR and SAA based risk approximations. As shown in the next section, this directly translates to lower constraint violations even while using very few number of rollouts and even fewer constraint evaluations.
Connections to Works on RKHS Embedding: Our formulation builds upon the RKHS embedding approach presented in [6], [17] for functions of multiple random variables. These works also suggest using the concept of reduced-set to maximize the expressive capacity of the embeddings. However [6], [17] relies on randomly selecting a subset of the samples as the reduced-set. In contrast, our bi-level optimization (11a)-(LABEL:bi_level_inner_red_cost) provides a much more principled approach. Our proposed work also extends the MMD-based risk surrogate presented in [15], [16] to the case of stochastic dynamics.
V Validation and Benchmarking
The objective of this section is to compare trajectory optimization with $r_{MMD}^{emp}$ as the risk cost vis-a-vis other popular alternatives. While doing so, we also expose the inner workings of our MMD-based approach.
V-A Implementation Details
We implemented optimization (11a)-(LABEL:bi_level_inner_red_cost) and Alg.1 in Python using Jax [25] as the GPU-accelerated linear algebra back-end. Our constraint function $h(\mathbf{x})$ has two parts. For the first part, we check if every lateral position $d_{k}$ is within the lane bounds. The second part computes the distance of the ego-vehicle with the neighboring vehicles, by modeling each of them as ellipsoids. We handle multiple obstacles by simply taking the worst-case collision distance over all the obstacles [8]. To ensure the reproducibility of our benchmarking, we consider only static obstacles in our comparative analysis and we assume that their positions are known. But in the accompanying video, we show results with dynamic obstacles as well.
The state cost has the following form
$c(\mathbf{x})=\sum_{k}(v_{k}-v_{d})^{2}+|(d_{k}-d_{1})||(d_{k}-d_{2})|+(\ddot{s}_{k}^{2}+\ddot{d}_{k}^{2}),$ (14)
where $v_{d}$ is the desired forward velocity and $d_{1},d_{2}$ are two lane center-lines which the ego-vehicle can choose to follow at any given time. The last two terms penalize value in the second-order position derivatives.
Figure 3: (3 Static obstacles) Ground-truth collision-rate achieved by minimum $r_{MMD}^{emp}$ and $r_{CVaR}^{emp}$ trajectories (see (17)). We ensured that the respective optimal trajectories achieves $r_{MMD}^{emp}=0$, $r_{CVaR}^{emp}=0$. Thus, this result showcases the finite sample effectiveness of our MMD-based risk cost at low-sample regimes. We use the following values of the positive constants in 15a, 15b, 16a, 16b: Low Gaussian noise: $c_{a,1}=0.1,c_{a,2}=0.001,c_{\theta,1}=0.1,c_{\theta,2}=0.001$. High Gaussian noise: $c_{a,1}=0.15,c_{a,2}=0.001,c_{\theta,1}=0.15,c_{\theta,2}=0.001$. Low Beta noise: $c_{a,1}=0.1,c_{a,2}=0.001,c_{\theta,1}=0.001,c_{\theta,2}=0.001$. High Beta noise $c_{a,1}=0.15,c_{a,2}=0.001,c_{\theta,1}=0.0015,c_{\theta,2}=0.001$
Figure 4: (1 dynamic obstacle) Ground-truth collision-rate achieved by minimizing $r_{MMD}^{emp}$ and $r_{CVaR}^{emp}$ (see (17)). The optimizer achieved $r_{MMD}^{emp}=0$, $r_{CVaR}^{emp}=0$ in all the scenarios. Thus, this result showcases the finite sample effectiveness of our MMD-based risk cost at low-sample regimes. We use the following values of the positive constants in 15a, 15b, 16a, 16b: Low Gaussian noise: $c_{a,1}=0.1,c_{a,2}=0.001,c_{\theta,1}=0.1,c_{\theta,2}=0.001$. High Gaussian noise: $c_{a,1}=0.15,c_{a,2}=0.001,c_{\theta,1}=0.15,c_{\theta,2}=0.001$. Low Beta noise $c_{a,1}=0.1,c_{a,2}=0.001,c_{\theta,1}=0.005,c_{\theta,2}=0.001$. High Beta noise: $c_{a,1}=0.15,c_{a,2}=0.001,c_{\theta,1}=0.0075,c_{\theta,2}=0.001$
(a) $r_{MMD}^{emp}$
(b) $r_{CVaR}^{emp}$
(c) $r_{MMD}^{emp}$
(d) $r_{CVaR}^{emp}$
Figure 6: ( MPC simulations in Carla T5 (a)-(b) and T10 (c)-(d) town. The red line is the route/reference-path. We show trajectories (green line) resulting from minimizing $r_{MMD}^{emp}$ (left column) and $r_{CVaR}^{emp}$ (right column) for $N=2$. The small number of samples of constraint residuals provide a very unreliable estimate of $r_{CVAR}^{emp}$ that changes sharply between consecutive MPC steps. As a result, the ego-vehicle is often unable to find any safe trajectory. As in the case shown, $r_{CVaR}^{emp}$ driven vehicle gets stuck behind the obstacle and eventually collides. In sharp contrast, $r_{MMD}^{emp}$ is able to regularly provide a consistent risk estimate for guiding the ego-vehicle. For town T5, the control noise for this particular experiment followed the Beta distribution with $c_{a,1}=0.05,c_{a,2}=0.4,c_{\theta,1}=0.05,c_{\theta,2}=0.01$ and for town T10, the control noise followed the Gaussian distribution with $c_{a,1}=0.3,c_{a,2}=0.3,c_{\theta,1}=0.3,c_{\theta,2}=0.01$
V-A1 Benchmarking Environment
We perform benchmarking in trajectory-optimization and Model Predictive Control (MPC) setting. For the former, we randomly sample the initial states of the ego-vehicle, static obstacles and trajectories of dynamic obstacles, and use Alg.1 to compute an optimal trajectory. The latter involves constant re-planing using Alg.1 and the current state of the ego-vehicle.
For the benchmarking in MPC setting, we use high-fidelity simulator CARLA[12] to perform all our experiments. Each experiment requires the ego vehicle to maneuver from a start position to a goal position following a given route (reference path). An experiment is counted as successful if the ego vehicle makes a complete run without colliding with any of the obstacles. We perform experiments in two towns, namely, town 10(T10) and town 05(T5) . We consider a two-lane driving scenario with 8 and 10 static obstacles, respectively, in town 10 and town 05. We consider uncertainty in the dynamics of the ego vehicle, injected via perturbations of the nominal acceleration and steering control inputs.
The Gaussian noise distribution has the following form:
$\epsilon_{a,k}\sim|c_{a,1}a_{k}|\mathcal{N}(0,1)+c_{a,2}\mathcal{N}(0,1),$ (15a)
$\epsilon_{\theta,k}\sim|c_{\theta,1}\theta_{k}|\mathcal{N}(0,1)+c_{\theta,2}\mathcal{N}(0,1),$ (15b)
where $c_{a,i}$ and $c_{\theta,i}$ are positive constants. As can be seen, the noise have a part that depends on the magnitude of the control and a part that is constant. We create different noise settings by varying $c_{a,i}$ and $c_{\theta,i}$. We also consider a setting where the r.h.s of (15a), (15b) are replaced by Beta distribution with probability density function $g(x;a,b)\propto x^{a-1}(1-x)^{b-1}$ where $a,b$ are control dependent parameters. The Beta noise distribution has the following form:
$\epsilon_{a,k}\sim c_{a,1}\mathcal{B}(a,b)+c_{a,2}\mathcal{N}(0,1),$ (16a)
$\epsilon_{\theta,k}\sim c_{\theta,1}\mathcal{B}(a,b)+c_{\theta,2}\mathcal{N}(0,1),$ (16b)
$a=\{2|a_{k}|,2|\theta_{k}|\},b=\{5|a_{k}|,5|\theta_{k}|\}$ (16c)
V-A2 Baselines
We consider two baselines.
Derterministic (DET): This is a noise ignorant approach. We use [18] as the planner as its structure is very similar to Alg.1.
CVaR based Approach: This baseline is obtained by replacing $r_{MMD}^{emp}$ with a CVaR based risk defined in the following manner, in Alg.1.
$r(\mathbf{x})\approx r_{CVaR}^{emp}=CVaR^{emp}(\overline{h}(\mathbf{x}))$ (17)
where $CVaR^{emp}$ is the empirical CVaR estimate computed using the constraint residual samples $\overline{h}({{}^{ij}}\mathbf{x})$ along the state-trajectory rollouts following the approach of [1]. Since, the underlying trajectory optimizer is same for minimizing both our proposed risk surrogate $r_{MMD}^{emp}$ and the baseline $r_{CVaR}^{emp}$, our benchmarking is designed to minimize optimizer bias and study the finite-sample performance of each risk estimate.
V-A3 Metrics
We use the following 3 different metrics to validate our results as well as compare against the baseline.
Collision percentage: In the trajectory optimization setting, we rollout the vehicle states by perturbing the optimal trajectory computed from Alg.1. We perform a large number of rollouts to estimate the ground-truth collision-rate. In the MPC setting, we calculate the number of collisions out of 50 experiments ran in each of town T10 and T05. Each MPC run, consists of around 2000 re-planning steps.
Lane constraints violation percentage:. We don’t consider any lane-constraints in the trajectory optimization setting. In the MPC setting, for each experiment we add the lateral lane violations(in metres) at each MPC step and subsequently divide it by the total arc length of the reference trajectory. We then take the average across all experiments.
Average and Maximum speed(m/s): For average speed, we add the ego speed at each MPC step for each experiment and then divide by the total number of steps in that particular experiment. Subsequently we take the average across all experiments. For maximum speed, we compute the highest speed achieved in each MPC experiment and subsequently average it across experiments.
Remark 1.
In all our experiments, we fix the number of state-trajectory samples $N$ along which we evaluate $h(\mathbf{x})$ to estimate the risk. We recall, that our approach involves first computing $N^{2}$ state-trajectory rollouts and then choosing $N$ out of those for evaluating $h(\mathbf{x})$. This in turn limits the number of collision-checks that we need to perform. For CVaR baselines, the down-sampling process is not relevant and we compute only $N$ state-trajectory rollouts and evaluate $h(\mathbf{x})$ over them.
V-B Benchmarking in Trajectory Optimization Setting
We created two sets of benchmarks with static and dynamic obstacles. For the former, we created 200 random scenarios and use two different noise models to perturb nominal dynamics. For the later, we had a single dynamic obstacle moving along a pre-defined trajectory and here again, we had two different noise models. The qualtiative trajectories in some of these scenarios are presented in the accompanying video. We use Alg.1 to compute minimum $r_{MMD}^{emp}$ and $r_{CVaR}^{emp}$ trajectories in both static and dynamic obstacle benchmarks. Importantly, we ensured that the optimal trajectory obtained for each scenario in both the benchmarks had $r_{MMD}^{emp}=0$ and $r_{CVaR}^{emp}=0$. We then performed sampling around the optimal trajectory to compute the ground-truth collision-rate. To keep the analysis simple, this benchmarking did not enforce any risk cost associated with the lane violations.
The results are summarized in Fig.3,4 for different $N$. As can be seen that although both $r_{MMD}^{emp}$ and $r_{CVaR}^{emp}$ are zero, the real collision-rates are very different. Our $r_{MMD}^{emp}$ results in lower collision-rate than $r_{CVaR}^{emp}$. For example, for low Gaussian noise and $N=4$, $r_{MMD}^{emp}$ trajectories provide almost two-times lower collision-rate than $r_{CVaR}^{emp}$ trajectories. Moreover, the collision-rate along $r_{MMD}^{emp}$ optimal trajectories shrink faster than they do along $r_{CVaR}^{emp}$ optimal trajectories. This result showcases the effectiveness of finite-sample estimate of our risk cost based on MMD vis-a-vis CVaR. This superior performance of $r_{MMD}^{emp}$ can be attributed to the fact the RKHS embedding is particularly effective in capturing the true distribution of arbitrary function of random variables with only a handful of samples (see Fig.1 in [6]). Our optimal reduced-set method (11a)-(LABEL:bi_level_inner_red_cost) further supercharges the capability of the RKHS embedding.
V-C Benchmarking in MPC Setting using CARLA
We now evaluate the efficacy of $r_{MMD}^{emp}$ in a MPC setting where constant re-planning is done based on the current feedback of ego and the neighbouring vehicle state. In this benchmarking, we enforce a risk on the lane constraints as well. We also introduced Gaussian noise in the initial state.
Fig.6 presents a typical performance observed with $r_{MMD}^{emp}$ and $r_{CVaR}^{emp}$ under Gaussian noise distribution. Due to extremely small $N$, the variance in $r_{CVaR}^{emp}$ estimation between consecutive MPC steps is large. As a result, the ego-vehicle often gets stuck behind the static obstacles and eventually collides. In sharp contrast, $r_{MMD}^{emp}$ estimates are consistently successful in guiding the ego-vehicle without collision. We demonstrate several such-scenarios in the supplementary video.
Table I presents the quantitative benchmarking in the MPC setting. We observed that re-planning can counter some of the effects of noise. Thus both $r_{MMD}^{emp}$ and $r_{CVaR}^{emp}$ approaches show improvement over the performance we observed in a stand-alone planning setting of the previous sub-section. Nevertheless, $r_{MMD}^{emp}$ still outperformed $r_{CVaR}^{emp}$ in terms of collision-rate and achieved max and average speed. The lane violations of CVaR baseline was lower. However, this was because it had shorter runs due to collisions in several experiments. Table I also presents the results of the deterministic baseline. As can be seen, a noise ignorant approach is overconfident and moves too fast and consequently did not succeed in any of the simulation runs.
Table II repeats the benchmarking for higher noise level. Here, we only present results for our MMD-based approach and CVaR baseline. At higher noise, the performance (collision-rate, speed) of our $r_{MMD}^{emp}$ driven approach deteriorates a little, as compared to Table I. But, the degradation is far less than that observed for trajectories resulting from minimizing $r_{CVaR}^{emp}$.
TABLE I: $N=2$, Rollout horizon 40, 50 experiments
Gaussian $c_{a,1}=c_{\theta,1}=0.3$, Beta $c_{a,1}=c_{\theta,1}=0.01$
$c_{a,2}=0.3,c_{\theta,2}=0.01$, Gaussian noise in the initial state.
Method/Town % Collisions (Gaussian/Beta) % Lane constr. (Gaussian/Beta) Avg. Speed(m/s) (Gaussian/Beta) Max. Speed(m/s) (Gaussian/Beta)
$r_{MMD}^{emp}$/T5 0/0 2.39/1.18 2.59/2.68 3.69/5.08
$r_{CVaR}^{emp}$/T5 7.69/0 2.25/1.08 2.12/2.49 3.04/4.61
DET/T5 100/100 0.03/0 5.25/5.27 8.72/8.15
$r_{MMD}^{emp}$/T10 0/0 0.47/0.99 3.42/3.63 5.94/5.5
$r_{CVaR}^{emp}$/T10 16.67/0 0.07/0.55 2.8/3.02 5.19/5.17
DET/T10 100/100 1.01/0.48 5.8/5.93 10.5/9.61
TABLE II: $N=2$, Rollout horizon 40, 50 experiments
Gaussian $c_{a,1}=c_{\theta,1}=0.3$, Beta $c_{a,1}=c_{\theta,1}=0.05$
$c_{a,2}=0.4,c_{\theta,2}=0.01$, Gaussian noise in the initial state.
Method/Town % Collisions (Gaussian/Beta) % Lane constr. (Gaussian/Beta) Avg. Speed(m/s) (Gaussian/Beta) Max. Speed(m/s) (Gaussian/Beta)
$r_{MMD}^{emp}$/T5 0/0 2.7/3.5 2.59/2.02 4.06/4.07
$r_{CVaR}^{emp}$/T5 15/45 2.14/2.46 2.08/1.53 3.45/3.4
$r_{MMD}^{emp}$/T10 4/0 0.5/5.11 3.56/2 7.01/3.73
$r_{CVaR}^{emp}$/T10 17/4 0.1/3.23 2.73/1.56 4.55/4.07
VI Conclusions and Future Work
Estimating state-dependent risk through forward simulation is a popular approach in motion planning. But computing rollouts can be expensive. More critically, evaluating constraints over the rollouts could be even more computationally prohibitive, for example, due to expensive collision checks. Thus, we presented a first such principled approach for reducing the number of constraint evaluations needed to estimate risk. Furthermore, our work can produce reliable results with a very few number of rollouts of system dynamics. Specifically, our work leveraged the solid mathematical foundations of RKHS embedding to propose a risk surrogate whose finite sample effectiveness can be improved by solving a bi-level optimization. We performed extensive simulations in both stand-alone trajectory optimization as well as MPC setting with a strong baseline based on CVaR. We showed that our MMD-based approach outperforms it in collision-rate and achieved average and maximum speed.
Currently the entire computations of Alg.1 can run between 5-10 Hz on a RTX 3090 desktop. We believe this performance can be improved by learning good warm-start for Alg.1. We are also looking to replace optimization (11a)-(LABEL:bi_level_inner_red_cost) with a neural network that can directly predict the optimal reduced-set. Our future works also seek to extend the formulation to different robotics systems like quadrotors, manipulators, etc.